OpenStudy (anonymous):

there are 1500 students in the school based on the sample proportion, how many students in the school would be expected to do household chores for at least two hours every week? Time (in hours) 0-1 2-3 4-5 6-7 Number of students 32 64 3 1 A.480 B.960 C.1.020 D.1.440

OpenStudy (anonymous):

PlZ help

OpenStudy (anonymous):

i don't understand this question

OpenStudy (kropot72):

Number of students in the sample = 32 + 64 + 3 + 1 = 100 Number of students in the sample who do chores at least 2 hrs/week = 64 + 3 + 1 = 68 Proportion of students in the sample who do chores at least 2 hrs/week = 68/100 Expected number of students in the school who do chores at least 2 hrs/week is given by: \[\large Expected\ number=\frac{1500\times68}{100}=you\ can\ calculate\]